OpenStudy (anonymous):

The mathematics department of a college has 12 male professors, 10 female professors, 11 male teaching assistants, and 6 female teaching assistants. If a person is selected at random from the group, find the probability that the selected person is a teaching assistant or a female.

OpenStudy (anonymous):

counting for this one

OpenStudy (anonymous):

\(10+12+11+6=39\) that will be your denominator, there are \(10+6=16\) females and another 11 teaching assistants, for a total of \(16+11=27\) favourable to your question